# Asymptotics for conformal inference

### v1

@@ 4.2 Main results
When $P_{0}=P_{\tiny\mbox{cal}}$ (no distribution shift), the following result holds for the regular Bates et al., (2023) procedure.
@@
Under Assumption 2 with $P_{0}=P_{\tiny\mbox{cal}}$, Assumption 3 and and Assumption 5, let us consider $\mathcal{R}_{\alpha}$ the BH procedure at level $\alpha$ applied to the conformal $p$-values (4), with a level $\alpha>[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$. If $n/(n+m)\rightarrow\sigma^{2}\in[0,1]$ and $\pi_{0}(m)\rightarrow\pi_{0}\in(0,1)$, we have
$\sqrt{\tau_{n,m}}\left(\mathrm{FDP}_{m}^{(n)}(\mathcal{R}_{\alpha})-\pi_{0}\alpha\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\left(0,\alpha^{2}\pi_{0}\left[\sigma^{2}+(1-\sigma^{2})\pi_{0}\right]\frac{1-\mathcal{T}_{\alpha}}{\mathcal{T}_{\alpha}}\right);$ (25)
$\sqrt{\tau_{n,m}}\left(\mathrm{TDP}_{m}^{(n)}\left(\mathcal{R}_{\alpha}\right)-G(\mathcal{T}^{*})\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\left(0,\frac{\Sigma_{\alpha}}{(\alpha^{-1}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha}))^{2}}\right),$ (26)
with $\Sigma_{\alpha}={{G^{\prime}(\mathcal{T}_{\alpha})^{2}\mathcal{T}_{\alpha}(1-\mathcal{T}_{\alpha})\left[\pi_{0}\sigma^{2}+(1-\sigma^{2}){\alpha^{-2}}\right]}}{}+{\left[\alpha^{-1}-\pi_{0}\right]^{2}\left(1-\pi_{0}\right)^{-1}G(\mathcal{T}_{\alpha})(1-G(\mathcal{T}_{\alpha}))\sigma^{2}}{}.$
Theorem 4.1 is a direct corollary of Proposition 5.3, itself proved in Section 5.2.3. It relies on the pipeline introduced by Neuvial, (2008), which consists in first deriving a functional central limit theorem for the e.c.d.f. of the $p$-values and then to use the functional delta method (van der Vaart,, 1998).
@@
$\sqrt{m}\left(\mathrm{FDP}_{m}^{(n)}(\mathcal{R}_{\alpha})-\pi_{0}\alpha\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\left(0,\alpha^{2}\pi_{0}\left[1+(\sigma^{-2}-1)\pi_{0}\right]\frac{1-\mathcal{T}_{\alpha}}{\mathcal{T}_{\alpha}}\right).$
When $\sigma=1$ (that is $n/m\to\infty$), we note that the FDP convergence is the same as when the $p$-values are independent (Neuvial,, 2008). However, when $n/m$ is bounded, the asymptotic variance is affected by the dependence and gets larger when $\sigma$ decreases (that is, $n/m$ decreases). This is coherent with results on FDP convergence in literature dealing with dependency structure: the (positive) dependence is shown to increase the dispersion of the FDP, see Delattre and Roquain, (2011, 2016) among others.
@@
$\sqrt{\tau_{n,m}}\left(\mathrm{FDP}_{m}^{(n)}(\mathcal{R}^{w^{*}}_{\alpha})-\pi_{0}\alpha\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\left(0,\alpha^{2}\pi_{0}\Xi^{{w^{*}}}_{\alpha}\frac{1-\mathcal{T}^{{w^{*}}}_{\alpha}}{\mathcal{T}^{{w^{*}}}_{\alpha}}\right);$ (27)
$\sqrt{\tau_{n,m}}\left(\mathrm{TDP}_{m}^{(n)}(\mathcal{R}^{w^{*}}_{\alpha})-{G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\left(0,\frac{\Sigma^{{w^{*}}}_{\alpha}}{(\alpha^{-1}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha}))^{2}}\right),$ (28)
@@
$\Sigma^{{w^{*}}}_{\alpha}$ $={{\left({(G^{w^{*}})}^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})\right)^{2}\mathcal{T}^{{w^{*}}}_{\alpha}(1-\mathcal{T}^{{w^{*}}}_{\alpha})\pi_{0}\sigma^{2}}}{}$
@@
$\ +{\left({\rho^{w^{*}}}{(G^{w^{*}})}^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha}){\alpha^{-1}}\right)^{2}\left[\mathcal{T}^{{w^{*}}}_{\alpha}-I^{w^{*}}(\mathcal{T}^{{w^{*}}}_{\alpha})^{2}\right](1-\sigma^{2})}{}$
$\ +{\left[\alpha^{-1}-\pi_{0}\right]^{2}\left(1-\pi_{0}\right)^{-1}{G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})(1-{G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha}))\sigma^{2}}{}.$
@@
Theorem 4.2 is proved in Section 5.2.4. The Bates et al., (2023) procedure with (similar) weighted conformal $p$-values has been studied in Jin and Candès, (2023) and they obtained a convergence of FDR and FDP to quantities analogue to $\pi_{0}\alpha$ and ${G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})$, respectively. By contrast, Theorem 4.2 provides the full asymptotic distribution.
@@
While Theorem 4.2 focuses on the oracle weight function, the non-oracle case is deferred to Section D, see Theorem D.2 .

### v3

@@ 4.2 Main results
When $P_{0}=P_{\tiny\mbox{cal}}$ (no distribution shift), the following result provides the convergence behavior of the novelty detection procedure of Bates et al. (2023), that is, of $\mathcal{R}_{\alpha}$ given in Section 2.2.
@@
Under Assumption 2 with $P_{0}=P_{\tiny\mbox{cal}}$, Assumption 3 and Assumption 5, let us consider $\mathcal{R}_{\alpha}$ the BH procedure at level $\alpha$ applied to the conformal $p$-values (4), with a level $\alpha>[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$. If $n/(n+m)\rightarrow\sigma^{2}\in[0,1]$ and $\pi_{0}(m)\rightarrow\pi_{0}\in(0,1)$, we have
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}(\mathcal{R}_{\alpha})-\pi_{0}\alpha}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\mathopen{}\mathclose{{\left(0,\alpha^{2}\pi_{0}\mathopen{}\mathclose{{\left[\sigma^{2}+(1-\sigma^{2})\pi_{0}}}\right]\frac{1-\mathcal{T}_{\alpha}}{\mathcal{T}_{\alpha}}}}\right);$ (26)
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{TDP}_{m}^{(n)}\mathopen{}\mathclose{{\left(\mathcal{R}_{\alpha}}}\right)-G(\mathcal{T}_{\alpha})}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\mathopen{}\mathclose{{\left(0,\frac{\Sigma_{\alpha}}{(\alpha^{-1}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha}))^{2}}}}\right),$ (27)
with $\Sigma_{\alpha}={{G^{\prime}(\mathcal{T}_{\alpha})^{2}\mathcal{T}_{\alpha}(1-\mathcal{T}_{\alpha})\mathopen{}\mathclose{{\left[\pi_{0}\sigma^{2}+(1-\sigma^{2}){\alpha^{-2}}}}\right]}}{}+{\mathopen{}\mathclose{{\left[\alpha^{-1}-\pi_{0}}}\right]^{2}\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)^{-1}G(\mathcal{T}_{\alpha})(1-G(\mathcal{T}_{\alpha}))\sigma^{2}}{}.$
Theorem 4.1, proved in Section 5.2.3, relies on the pipeline introduced by Neuvial (2008), which consists in first deriving a functional central limit theorem for the e.c.d.f. of the $p$-values and then using the functional delta method (van der Vaart, 1998).
@@
$\sqrt{m}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}(\mathcal{R}_{\alpha})-\pi_{0}\alpha}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\mathopen{}\mathclose{{\left(0,\alpha^{2}\pi_{0}\mathopen{}\mathclose{{\left[1+(\sigma^{-2}-1)\pi_{0}}}\right]\frac{1-\mathcal{T}_{\alpha}}{\mathcal{T}_{\alpha}}}}\right).$
When $\sigma=1$ (that is $n/m\to\infty$), we note that the FDP convergence is the same as when the $p$-values are independent (Neuvial, 2008). However, when $n/m$ is bounded, the asymptotic variance is affected by the dependence and gets larger when $\sigma$ decreases. In addition, and maybe more importantly, the convergence rate is heavily affected in the under-calibrated setting when $\sigma^{2}=0$, with a convergence rate $\sqrt{\tau_{n,m}}\ll\sqrt{m}$. This phenomenon is illustrated in Figure 3 for $m=10^{3}$ and various $n$ in a particular setting, with corresponding convergence rates reported in Table 2. The three different regimes from the prediction setting displayed in Table 1 also appear in this novelty detection setting.
Figure 3: Asymptotic approximation of the distribution of the $\mathrm{FDP}$ of the Benjamini-Hochberg procedure with conformal $p$-values given by Theorem 4.1 with $m=10^{3}$ and several choice of $n$ (plain lines) at level $1-\alpha=0.8$. The case of independent theoretical $p$-values (dashed line) is displayed for comparison. The null distribution and the calibration distribution is $\mathcal{N}(0,1)$ while the distribution of the alternative is $\mathcal{N}(3,1)$, with a null proportion $\pi_{0}$ equal to $70\%$.
Calibration sample size: $n$ Rate of convergence: $\sqrt{\tau_{n,m}}$ Independence coefficient: $\sigma^{2}$
Conformal $\sqrt{m}$ $m^{1/4}$ 0
$\frac{m}{\log(m)}$ $\sqrt{\frac{m}{\log(m)}}$ 0
$\frac{m}{2}$ $\sqrt{\frac{m}{3}}$ $\frac{1}{3}$
$m$ $\sqrt{\frac{m}{2}}$ $\frac{1}{2}$
$2m$ $\sqrt{\frac{2m}{3}}$ $\frac{2}{3}$
$\gamma m$, $\gamma>0$ $\sqrt{\frac{\gamma}{\gamma+1}m}$ $\frac{\gamma}{\gamma+1}$
$\log(m) m\sqrt{m}$ 11
$^{3/2} m\sqrt{m}$ 11
$^{2} m\sqrt{m}$ 11
Independent $\infty$ $\sqrt{m}$ 11
Table 2: Rates $\sqrt{\tau_{n,m}}$ and $\sigma^{2}$ in Theorems 3.23.34.1 and 4.2, with different choices of $n$ and $m$. Except the case $n=\gamma m$, they all correspond to the different regimes displayed in Figure 3. The case with independent $p$-values is presented for comparison.
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}(\mathcal{R}^{w^{*}}_{\alpha})-\pi_{0}\alpha}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\mathopen{}\mathclose{{\left(0,\alpha^{2}\pi_{0}\Xi^{{w^{*}}}_{\alpha}\frac{1-\mathcal{T}^{{w^{*}}}_{\alpha}}{\mathcal{T}^{{w^{*}}}_{\alpha}}}}\right);$ (28)
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{TDP}_{m}^{(n)}(\mathcal{R}^{w^{*}}_{\alpha})-{G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathcal{N}\mathopen{}\mathclose{{\left(0,\frac{\Sigma^{{w^{*}}}_{\alpha}}{(\alpha^{-1}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha}))^{2}}}}\right),$ (29)
@@
$\Sigma^{{w^{*}}}_{\alpha}$ $={{\mathopen{}\mathclose{{\left({(G^{w^{*}})}^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})}}\right)^{2}\mathcal{T}^{{w^{*}}}_{\alpha}(1-\mathcal{T}^{{w^{*}}}_{\alpha})\pi_{0}\sigma^{2}}}{}$
@@
$\ +{\mathopen{}\mathclose{{\left({\rho^{w^{*}}}{(G^{w^{*}})}^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha}){\alpha^{-1}}}}\right)^{2}\mathopen{}\mathclose{{\left[\mathcal{T}^{{w^{*}}}_{\alpha}-I^{w^{*}}(\mathcal{T}^{{w^{*}}}_{\alpha})^{2}}}\right](1-\sigma^{2})}{}$
$\ +{\mathopen{}\mathclose{{\left[\alpha^{-1}-\pi_{0}}}\right]^{2}\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)^{-1}{G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})(1-{G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha}))\sigma^{2}}{}.$
@@
Theorem 4.2 is proved in Section 5.2.4. The weighted conformal BH procedure with (similar) weighted conformal $p$-values has been studied in Jin and Candès (2023) and they obtained a convergence of FDR and TDR (i.e. $\mathbb{E}[\mathrm{FDP}_{m}^{(n)}(\mathcal{R}^{w^{*}}_{\alpha})]$ and $\mathbb{E}[\mathrm{TDP}_{m}^{(n)}(\mathcal{R}^{w^{*}}_{\alpha})]$) to quantities analogue to $\pi_{0}\alpha$ and ${G^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})$, respectively. By contrast, Theorem 4.2 provides the full asymptotic distribution.
@@
While Theorem 4.2 focuses on the oracle weight function, the non-oracle case is deferred to Section D (see Theorem D.2) .
The limiting distribution in Theorems 4.1 and 4.2 are similar to the previous ones in the literature (Neuvial, 2008; Delattre and Roquain, 2011, 2016; Kluger and Owen, 2024) which is well expected since all these works study the asymptotic normality of the $\mathrm{FDP}$ and the $\mathrm{TDP}$ of the $\mathrm{BH}$ procedure. However, the novelty is in the setting: we do not have access to the theoretical and uniform $p$-values under the null, but only to the conformal $p$-values. Conformal $p$-values are by nature dependent (and PRDS for the unweighted ones (Bates et al., 2023)) even with independent scores, while the dependency structures on the theoretical $p$-values in Delattre and Roquain (2011, 2016); Kluger and Owen (2024) come from the dependency of the gaussian test statistics. Furthermore conformal $p$-values are approximations of the theoretical ones, hence the study of the double asymptotic in both the nimber of test $m$ (which is the only one studied in the literature) and the size $n$ of the calibration set is crucial and provided here. Equivalent of the rate of convergence $\sqrt{\tau_{n,m}}$ and the limit $\sigma^{2}$ in Theorems 3.2 and 4.1 and their weighted version Theorems 3.3 and 4.2 for different choice of $n$ depending on $m$.
4.3 Applications
Theorem 4.1 has several possible applications. A first application is to obtain an asymptotic confidence region of the $\mathrm{FDP}$ of the BH procedure in the conformal case.
Corollary 4.3.
Let $\delta>0$. Under Assumption 2 with $P_{0}=P_{\tiny\mbox{cal}}$ and Assumptions 3 and 5, let us consider $\mathcal{R}_{\alpha}$ the BH procedure at level $\alpha$ applied to $m$ conformal $p$-values (4) with a calibration sample of size $n$, with a level $\alpha>[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$. Denote $\widehat{\mathcal{T}}_{\alpha}=\alpha\mathopen{}\mathclose{{\left|\mathcal{R}_{\alpha}}}\right\rvert/m$ the threshold of $\mathcal{R}_{\alpha}$. Define the upper bound:
$\widehat{\lambda}_{n,m}(\delta,\alpha,\pi_{0})=\alpha\sqrt{\pi_{0}\mathopen{}\mathclose{{\left[\frac{1}{m}+\frac{1}{n}\pi_{0}}}\right]\frac{1-\widehat{\mathcal{T}}_{\alpha}}{\widehat{\mathcal{T}}_{\alpha}}}\Phi^{-1}(1-\delta),$
with $\Phi^{-1}$ the quantile function of the standard gaussian distribution. Then if $n/(n+m)\rightarrow\sigma^{2}\in[0,1]$ and $\pi_{0}(m)\rightarrow\pi_{0}\in(0,1)$, we have
$\mathbb{P}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}(\mathcal{R}_{\alpha})\leqslant\pi_{0}\alpha+\widehat{\lambda}_{n,m}(\delta,\alpha,\pi_{0})}}\right)\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}1-\delta.$
[…]
${\widehat{\pi}}_{0}^{{\tiny{\mbox{Storey}}}}(m,n)=\frac{1+\sum_{i=1}^{m}{\mathds{1}_{{p_{i}^{(m)}}\geqslant\lambda}}}{m(1-\lambda)},$ (31)
where $\lambda\in(0,1)$ is a fixed parameter. Indeed, Applying Theorem 5.3 (or more precisely arguments in its proof, because the convergence of all empirical processes holds almost surely by the Glivenko-Cantelli theorem and by continuity), we have
${\widehat{\pi}}_{0}^{{\tiny{\mbox{Storey}}}}(n,m){{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}}\pi_{0}+(1-\pi_{0})\frac{1-G(\lambda)}{1-\lambda}\geqslant\pi_{0}\ a.s.$
Note that using the Storey estimator in the context of conformal $p$-values has been already considered in the literature (e.g., Bates et al. (2023); Marandon et al. (2024)), but not for deriving an FDP bound.
Another application of Theorem 4.1 is to use the asymptotic distribution in (27) to design the score function $S:\mathcal{Z}\rightarrow\mathbb{R}$, see Section 2.2. Recall that the latter is computed from an independent training sample $\mathcal{D}_{{\tiny\mbox{train}}}$. For simplicity, assume that $\mathcal{D}_{{\tiny\mbox{train}}}=\mathopen{}\mathclose{{\left\{(Z_{i},h_{i})\in\mathcal{Z}\times\mathopen{}\mathclose{{\left\{0,1}}\right\},i\in\llbracket n_{{\tiny\mbox{train}}}\rrbracket}}\right\}$ is a sample containing both labeled data under the null ($h_{i}=0$) and under the alternative ($h_{i}=1$), so that the score function can typically be trained as the best classifier of null versus alternative (Marandon et al., 2024), or as the likelihood ratio of null and alternative “conformal $p$-values” leading to some pruning step (Liang, Sesia and Sun, 2024). The asymptotic distribution in (27) delivers additional insight, and allows to use the asymptotic distribution to design a score function in a more $\mathrm{TDP}$ (or broadly “multiple testing”) oriented way. Namely, let us assume that the score function $S$ lies in a parametric class $\mathopen{}\mathclose{{\left\{S_{\theta},\theta\in\Theta}}\right\}$ of score functions (for example a classifier based on a neural network). The classical machine learning approach is to seek for the best two-class classifier by minimising the empirical risk $\theta\in\Theta\mapsto n_{{\tiny\mbox{train}}}^{-1}\sum_{i\in\llbracket n_{{\tiny\mbox{train}}}\rrbracket}{\mathds{1}_{S_{\theta}(Z_{i})\neq h_{i}}}$ which is an empirical substitute of $\mathbb{P}(S_{\theta}(Z_{1})\neq h_{1})$ computed on $\mathcal{D}_{{\tiny\mbox{train}}}$. Alternatively, let us consider the situation where the user seeks to have the “majority” of the $\mathrm{TDP}$ above a specific power level $\beta\in(0,1)$, while conserving the finite-sample $\mathrm{FDR}$ control given by the conformal BH procedure. In this case, the task is to choose $\theta$ that maximises $\mathbb{P}(\mathrm{TDP}_{m}^{(n)}>\beta)\approx 1-\Phi(\sqrt{\tau_{n,m}}L_{\alpha}(\beta))$ with $L_{\alpha}(\beta)=(\beta-G(\mathcal{T}_{\alpha}))(\alpha^{-1}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha}))\Sigma_{\alpha}^{-1/2}$ by using Theorem 4.1. A natural empirical substitution leads to minimise
$\theta\in\Theta\mapsto(\beta-\widehat{G}_{\theta})({\widehat{\mathcal{T}}}_{\theta,\alpha})(\alpha^{-1}-{\widehat{G}}_{\theta,{\tiny\mbox{mixt}}}^{\prime}({\widehat{\mathcal{T}}}_{\theta,\alpha}))\widehat{\Sigma}_{\theta,\alpha}^{-1/2},$ (32)
where ${\widehat{G}}_{\theta}$ is a differentiable estimate based on $\mathcal{D}_{{\tiny\mbox{train}}}$ (e.g. a kernel estimator or a local polynomial estimator) of $G$ (14) when the score function is $S_{\theta}$ (i.e. of $1-F_{\theta,\tiny\mbox{test}}\circ F_{\theta,\tiny\mbox{cal}}^{-1}(1-\cdot)$ where for all $t\in\mathbb{R}$, $F_{\theta,\tiny\mbox{test}}(t)$ (resp. $F_{\theta,\tiny\mbox{cal}}(t)$) is equal to $\mathbb{P}(S_{\theta}(Z_{i})\leqslant t)$ with $i\in{\mathcal{H}}_{0}$ (resp. $i\in{\mathcal{H}}_{1}$)), its derivative is denoted ${\widehat{G}}^{\prime}_{\theta}$ and all the other quantities are plug-in estimates of $G_{{\tiny\mbox{mixt}}}$ (21), $\mathcal{T}_{\alpha}$ (46) and $\Sigma_{\alpha}$ defined in (27) with $G$ and $G^{\prime}$ replaced by ${\widehat{G}}_{\theta}$ and ${\widehat{G}}^{\prime}_{\theta}$, the function $G_{0}$ replaced by $I$ and $\pi_{0}$ replaced by $n_{{\tiny\mbox{train}}}^{-1}\sum_{i\in\llbracket n_{{\tiny\mbox{train}}}\rrbracket}(1-h_{i})$. More generally, other objective functions can be optimised by considering other $\mathrm{TDP}$ related quantities approximated by the asymptotic distribution (27) or by using versions of the empirical risk that are penalised by (32). Finally, a similar approach can be followed in the case of a distribution shift by applying Theorem 4.2 (or its generalisation Theorem D.2). In this case, the oracle weight should also be seen as a function of $\theta$ (if the oracle weight can be known for every score function $S_{\theta}$, $\theta\in\Theta$) or must be chosen among a parameterised family of weight function $\mathopen{}\mathclose{{\left\{w_{\lambda,\theta},\ (\lambda,\theta)\in\Lambda\times\Theta}}\right\}$ (when only some information on the distribution shift is known) and the training sample should contain both clean and shifted data. This heuristic way to develop score (and weight) function with a $\mathrm{TDP}$ focus and not only a best “marginal” classification is left for future work.